# Understanding Transformers via $N$-gram Statistics

### v1

@@ D.3 TinyStories
Figure 10: Rule selection for a TinyStories heldout sequence using $\mathcal{R}^{\textrm{subgram}}_{7}$. Analogous to Figure 5 but with optimal rule chosen from $\mathcal{R}^{\textrm{subgram}}_{7}$ instead of $\mathcal{R}^{\textrm{all}}_{7}$.
Figure 11: Rule selection for a TinyStories heldout sequence using $\mathcal{R}^{\textrm{suffix}}_{7}$. Analogous to Figure 5 but with optimal rule chosen from $\mathcal{R}^{\textrm{suffix}}_{7}$ instead of $\mathcal{R}^{\textrm{all}}_{7}$.
D.4 Wikipedia
We present analogous results in Section 7 for a 1.4B model trained on Wikipedia. In Table 5 we present the analogue of Table 2 (except we go up to maximum context length $M=6$). To investigate sensitiy to the choice of distance measure on probability distributions, for this section, optimal rules are chosen using the $L^{\infty}$ metric
$d_{\infty}(p,q)=\max_{i}|p_{i}-q_{i}|$ (14)
instead of the variational distance.
The top1-accuracy when using optimal rules from $\mathcal{R}^{\textrm{all}}_{6}$ is 64.5%. As with TinyStories, we see significant gains in accuracy when we increase rule strength. Achieving the number 64.5% (versus the corresponding 74.0% number for TinyStories from Table 1), perhaps a surprisingly a high score, is the result of two competing factors: on the one-hand, Wikipedia has a much greater diversity of $N$-gram statistics (which makes prediction harder), while on the other hand, the training data has more $N$-grams for use by the rules. Note that our reference LLM (a 1.4B model) achieves 50.1% top-1 accuracy on the 100 validation stories and a train loss of around 2.1 nats at the end of training.
ruleset / $M$ 1 2 3 4 5 6
$\mathcal{R}^{\textrm{all}}_{M}$ 26.1 41.4 50.5 56.0 60.5 64.5
$\mathcal{R}^{\textrm{subgram}}_{M}$ 26.1 40.7 48.6 52.5 55.0 57.2
$\mathcal{R}^{\textrm{suffix}}_{M}$ 26.1 40.5 47.7 50.4 51.3 51.9
$\textrm{backoff}_{M}$ 26.1 38.6 43.2 43.3 42.6 42.5
Table 5: Approximation Strength for Wikipedia. We look at the average top1-accuracy between optimal rule and LLM predictions for rules of varying strength and maximum context length. We compute this average over each token prediction from 10 holdout Wikipedia sequences each consisting of 2048 tokens.
@@ D.4 Wikipedia
Figure 12: Rule selection for a Wikipedia heldout sequence. Analogous to Figure 5 but with optimal rule chosen from $\mathcal{R}^{\textrm{all}}_{6}$ and with variational distance replaced with the $L^{\infty}$ metric for measuring distances between probability distributions.
Figure 13: Rule selection for a Wikipedia heldout sequence. Analogous to Figure 5 but with optimal rule chosen from $\mathcal{R}^{\textrm{all}}_{6}$ and with variational distance replaced with the $L^{\infty}$ metric for measuring distances between probability distributions.

### v2

@@ D.4 Rule Selection
Figure 10: Rule selection for a TinyStories heldout sequence using $\mathcal{R}^{\textrm{subgram}}_{7}$. Analogous to Figure 5 but with optimal rule chosen from $\mathcal{R}^{\textrm{subgram}}_{7}$ instead of $\mathcal{R}^{\textrm{all}}_{7}$. Model size: 160M.
Figure 11: Rule selection for a TinyStories heldout sequence using $\mathcal{R}^{\textrm{suffix}}_{7}$. Analogous to Figure 5 but with optimal rule chosen from $\mathcal{R}^{\textrm{suffix}}_{7}$ instead of $\mathcal{R}^{\textrm{all}}_{7}$. Model size: 160M.
@@
Figure 12: Rule selection for a Wikipedia heldout sequence. Analogous to Figure 5 but with optimal rule chosen from $\mathcal{R}^{\textrm{all}}_{7}$ and with variational distance replaced with the $L^{\infty}$ metric for measuring distances between probability distributions. Model size: 1.4B.
Figure 13: Rule selection for a Wikipedia heldout sequence. Analogous to Figure 5 but with optimal rule chosen from $\mathcal{R}^{\textrm{all}}_{7}$ and with variational distance replaced with the $L^{\infty}$ metric for measuring distances between probability distributions. Model size: 1.4B.